Assume you are managing a food processing plant in Ethiopia. The demand function

for one of your product is given as Qd=50-2p. (8 marks)

a) Find the point price elasticity if price is 15 ETB? Is it elastic or inelastic?

b) How do you interpret the elasticity result?

c) In order to get more revenue what will be your recommendation. Is it to increase

price or decrease price? Why?

d) Describe at least four determinants of the price elasticity demand for the food

product?


1
Expert's answer
2021-06-13T17:39:36-0400

Solution:

a.). Point price elasticity of demand (PEd) = "\\frac{\\% \\triangle Qd}{\\% \\triangle P} = \\frac{ \\triangle Qd}{ \\triangle P} \\times \\frac{ P}{ Q} \n\n\u200b"

First derive Qd:

Qd = 50 – 2p

P = 15

Qd = 50 – 2(15)

Qd = 50 – 30 = 20

Qd = 20

Therefore PEd:

"\frac{ \triangle Qd}{ \triangle P} = -2"

Point price elasticity of demand (PEd) = "-2\times \frac{ 15}{ 20} = -1.5"


Point price elasticity of demand (PEd) = 1.5, and it is elastic.


b.). The PEd is greater than 1 which means the demand for the product is elastic. This can be interpreted as consumers being very sensitive to changes in price. A slight percentage increase in price will lead to a bigger percentage decrease in quantity demanded.

 

c.). In order to get more revenue, the best option will be to reduce the price of the product. When the demand is elastic at a given price level, the company should reduce its price since the percentage drop in price will result in an even larger percentage increase in the quantity sold, hence increasing total revenue.

 

d.). The four determinants of the price elasticity of demand for the food product include the following:

1.). Availability of close substitutes – The price elasticity of demand would be considered elastic when consumers can easily substitute the food product with another readily available product, which they regard as similar. On the other hand, if consumers cannot substitute the food product, then the product will experience inelastic demand.

 

2.). Income levels – The price elasticity of demand is lower for high-income earners compared to low-income earners. This is because high-income earners are not affected much by changes in prices compared to low-income earners who are highly affected by changes in prices of a product.

 

3.). Whether the good is a luxury or necessity – Products which highly-priced such as luxury, tend to be highly elastic since a fall in their prices will lead to an increase in demand compared to low-priced items such as essential goods which are inelastic and demand doesn’t change much with price changes.

 

4.). Price levels – Products that have high prices tend to be demand elastic, while products which have lower prices are demand inelastic since their demand doesn’t change much with changes in price.